# Controlled Erasure as a Building Block for Universal Thermodynamically-Robust Superconducting Computing

### v1

@@ VI Conclusion
Sec. V.2 established a connection between the device’s circuit parameters and the CE notation. Linking various CE protocols together by leveraging this relationship allowed for universal computations. Sec. V.4 demonstrated a NAND gate that employs only CEs. Due to its superconducting operation, the device performs CE-based universal computations at extremely low energy cost.
@@
Follow-on efforts explore the thermodynamic efficiency of this NAND gate, SPICE simulations of the protocol for varying circuit constructions, and investigations into implementing other useful and more complex logical operations.
@@ VII Acknowledgements
The authors thank Fabio Anza, Scott Habermehl, Jacob Hastings, and Gregory Wimsatt for helpful comments and discussions, as well as the Telluride Science Research Center for its hospitality during visits, and the participants of the Information Engines workshop there for their valuable feedback. This material is based upon work supported by, or in part by, U.S. Army Research Laboratory, U.S. Army Research Office grant W911NF-21-1-0048.
@@ Appendix A Control Erase Timescale Approximation Scheme
In this section, we detail an approximation scheme for the potential in Eq. (3) during the control erase (CE) protocol described in Sec III. Specifically, to represent the distribution of microstates being erased into the $01$ state, we approximate this situation as a particle rolling down a quadratic slope. From this, we calculate $t_{\mathrm{CE}}$—the time it takes for the particle to travel to the bottom of the slope.
First, note that in Fig. 3, the coordinate $\phi_{1}$ has negligible change in comparison to the change in $\phi_{2}$. With this in mind, we will consider the potential to be dependent only on $\phi_{2}$. Next, the height of the potential $U$ that the particle travels down is on the order of $U_{0}\sim 10^{-22}$: Essentially, the particle’s motion can be described in terms of $\phi_{2}$ without loss of generality. Then, there are two forces acting on the particle: (i) the force due to the potential $-\mathrm{d}U/\mathrm{d}\phi_{2}$, and (ii) the classical damping force $-\gamma\mathrm{d}\phi_{2}/\mathrm{d}t$. This means that the particle’s motion is described by Newtonian mechanics due to the influence of a dimension-full potential $U$ and damping coefficient $\gamma$, written as:
@@
First, we note that Eq. (3) is dimensionless, while Eq. (5) describes a particle’s motion under the influence of a potential with units of energy. To reconcile this, we must rewrite Eq. (5) in terms of dimensionless quantities. Ref. [35] detailed a strategy for converting dimension-full equations of motion—such as Eq. (5)—into dimensionless equations of motion. To do this, we write a dimensional quantity $z$ as $z=z^{\prime}z_{c}$ in which $z^{\prime}$ is the dimensionless representation of the quantity, while $z_{c}$ represents a dimensional scaling factor: Table 1 summarizes the circuit quantity scaling parameters $z_{c}$.
@@
Now, after appropriately making substitutions and simplifying, we write the dimensionless Eq. (5) as:
The approximation scheme of the potential will now be detailed. First, we write the dimension-full potential as $U(\phi_{2})=k\phi_{2}^{2}/2$, where $k$ now serves as the ‘spring constant’ of the potential. In dimensionless form, the potential reads: $U^{\prime}(\varphi_{2})=k^{\prime}\varphi_{2}^{2}/2$. This implies that $k_{c}=U_{c}/2\phi_{c}^{2}$, and furthermore $k^{\prime}=U^{\prime}/(\varphi_{2})^{2}$. Now, the particle’s motion is written as:
@@
which is analogous to the differential equation modelling a damped harmonic oscillator which oscillates at frequency $\omega^{2}=k^{\prime}\theta$ under damping coefficient $\lambda=\Lambda/2$. We deploy the ansatz: $\varphi_{2}(t)=A\mathrm{e}^{\alpha t}$ in which $\alpha=-\lambda\pm\sqrt{\lambda^{2}-\omega^{2}}=-\lambda\pm\Omega$. Depending on the value of $R$, $L$ and $C$, $\Omega>0$ or $\Omega<0$: These two situations will be categorized as ‘overdamped’ and ‘underdamped’, respectively, in the subsequent discussion.
@@ A.0.1 Overdamped
When $\Omega>0$, the solution to Eq. (7) is [41]:
@@
which describes an overdamped oscillator with initial conditions $A$ and $B$. To understand them, we first detail the consequences of the particle-like representation of the distribution of microstates. The average position of this distribution localizes to a position at any given time—represented by a particle in this case—while its average initial velocity will be zero. With this, the initial conditions are $(i)\ \varphi_{2}(t^{\prime}=0)=\varphi_{2}^{o}$, and $(ii)\ \mathrm{d}\varphi_{2}(t^{\prime}=0)/\mathrm{d}t^{\prime}=0$. Solving for these two initial conditions, and substituting the results back into Eq. (8) gives:
@@
Numerically solving Eq (9) for a final time $t^{\prime}=t_{f}^{\prime}$, and multiplying the result by $\sqrt{LC}$, yields $t_{\mathrm{CE}}=t_{f}^{\prime}\sqrt{LC}$.
@@ A.0.2 Underdamped
If $\Omega<0$, we must define $\widetilde{\omega}=\sqrt{\omega^{2}-\lambda^{2}}$ in order for the general solution to be written as [41]:
@@
Eq. (10) describes an underdamped harmonic oscillator for which the coefficient $D$ and phase $\theta$ are solved using initial conditions. Subsequently solving for them in the same manner as before, and substituting the results into Eq. (10), produces:
@@ Appendix B Dwell Time
First, $\omega_{p,l}$ is the plasma frequency in the $l$th flux direction. To understand it, we consider the dimension-full potential $U(\phi_{1},\phi_{2})$, perform a Taylor expansion around a minimum $(\phi_{1}=\phi_{1}^{*},\phi_{2}=\phi_{2}^{*})$, and only look to the second order terms whose partial derivatives are both either with respect to $\phi_{1}$ or $\phi_{2}$:
$U(\phi_{1}^{*},\phi_{2}^{*})$ $\sim\dfrac{1}{2}\dfrac{\partial^{2}U}{\partial\phi_{1}^{2}}\bigg{|}_{(\phi_{1}^{*},\phi_{2}^{*})}(\phi_{1}-\phi_{1}^{*})^{2}$
$\quad\quad+\dfrac{1}{2}\dfrac{\partial^{2}U}{\partial\phi_{2}^{2}}\bigg{|}_{(\phi_{1}^{*},\phi_{2}^{*})}(\phi_{2}-\phi_{2}^{*})^{2}~{}.$
We identify that a partial derivative taken with respect to $\phi_{l}$, and subsequently evaluated at the minimum, corresponds the spring constant $k_{l}$. From here, we take the corresponding plasma frequency to be $\omega_{l}=\sqrt{k_{l}/m}$ for which the dimension-full mass $m=C$.
Furthermore, $\Delta U_{b}$ indicates the barrier height between a local minima and a local maxima; $k_{B}$ is the Boltzmann constant; and $T$ is the circuit’s operating temperature. $\Gamma_{l}$ characterizes how many escape events are expected out of the minima per second. Finally, the dwell time is the reciprocal of Eq. (11): $t_{\mathrm{d},l}=1/\Gamma_{l}$. Explicitly, the plasma frequencies of interest for the dimensionless barriers $\Delta U_{01}^{\prime}$ ($\Delta U_{11}^{\prime}$ and $\Delta U_{10}^{\prime}$) correspond to the flux direction $l=1$ ($l=2$). Then, for values of $L$ and $T$ in Fig. 6, we select the minimum dwell time to calculate the robustness $t_{\mathrm{d}}/t_{\mathrm{CE}}$.
@@ Appendix C Control Erasures of Interest
Fig. 8 exhibits all control erasure protocols of interest. Successive executions of a subset of these protocols can lead to performing a NAND gate. Note that this is not an exhaustive illustration of possible control erasure protocols, but only represent those of this work’s interest. Table 2 shows the indices, signs and relationships of the relevant circuit parameters for a particular CE. In the notation of Sec. V.2, the non-zero barrier-control parameter is determined by $j$. Meanwhile, the magnitudes of the tilt parameters which offset the bias and barrier changes to the potential are related by $|\varphi_{i\mathrm{x}}|>|\varphi_{j\mathrm{x}}|$, respectively. By using the effective magnitudes of the circuit parameters detailed in Sec. V.3 and successively executing effective CE protocols, a robust NAND gate can be performed.
Figure 8: All control erasure protocols of interest for this work, displayed with arrow-based notation first shown in Fig. 3.
@@
Table 2: Tabulated summary of the relationship between experimental circuit parameters—including their indices and signs—and the CE notation introduced in Sec. III. Here, +()+\;(-)+ ( - ) corresponds to the sign of a particular circuit parameter, that being $>0\;(<0)$. Note that the non-zero barrier flux control parameter for a given CE is determined by $j$. Meanwhile, the magnitudes of the tilt parameters—which offset the bias (barrier) effect of $\mu$ $(\varphi_{j\mathrm{xdc}})$—are related by $|\varphi_{i\mathrm{x}}|>|\varphi_{j\mathrm{x}}|$, respectively.

### v2

@@ VI Conclusion
Section V.2 established a connection between the device’s circuit parameters and the CE notation. Linking various CE protocols together by leveraging this relationship allowed for universal computations. Section V.4 demonstrated a NAND gate that employs only CEs. Due to its superconducting operation, the device can perform CE-based universal computations at extremely low energy cost.
@@
Follow-on efforts explore the thermodynamic efficiency of this NAND gate, run SPICE simulations of the protocol for varying circuit constructions, and investigate implementing other useful and more complex logical operations.
@@ VII Acknowledgements
The authors thank Fabio Anza, Scott Habermehl, Jacob Hastings, Kuen Wai Tang, and Gregory Wimsatt for helpful comments and discussions, as well as the Telluride Science Research Center for its hospitality during visits, and the participants of the Information Engines workshop there for their valuable feedback. This material is based upon work supported by, or in part by, U.S. Army Research Laboratory, U.S. Army Research Office grant W911NF-21-1-0048.
@@ Appendix A Control Erase Timescale Approximation Scheme
The following details an approximation scheme for the potential in Eq. (3) during the control erase (CE) protocol described in Sec III. Specifically, to represent the distribution of microstates being erased into the $01$ state, we approximate this situation as a particle sliding down a quadratic slope. From this, we calculate $t_{\mathrm{CE}}$—the time it takes for the particle to travel to the bottom of the slope.
First, note that in Fig. 3, the coordinate $\phi_{1}$ has negligible change in comparison to the change in $\phi_{2}$. With this in mind, we will consider the potential to be dependent only on $\phi_{2}$. Next, the height of the potential $U$ that the particle travels down is on the order of $U_{0}\sim 10^{-22}$: Essentially, the particle’s motion can be described in terms of $\phi_{2}$ without loss of generality. Then, there are two forces acting on the particle: (i) the force due to the potential $-\mathrm{d}U/\mathrm{d}\phi_{2}$ and (ii) the classical damping force $-\gamma\mathrm{d}\phi_{2}/\mathrm{d}t$. This means that the particle’s motion is described by Newtonian mechanics due to the influence of a dimension-full potential $U$ and damping coefficient $\gamma$, written as:
@@
Observe that Eq. (3) is dimensionless, while Eq. (5) describes a particle’s motion under the influence of a potential with units of energy. To reconcile this, we must rewrite Eq. (5) in terms of dimensionless quantities. Reference [35] detailed a strategy for converting dimension-full equations of motion—such as Eq. (5)—into dimensionless equations of motion. To do this, we write a dimensional quantity $z$ as $z=z^{\prime}z_{c}$ in which $z^{\prime}$ is the dimensionless representation of the quantity, while $z_{c}$ represents a dimensional scaling factor. Table 1 summarizes the circuit quantity scaling parameters $z_{c}$.
Now, after appropriately making substitutions and further simplifying, we write the dimensionless Eq. (5) as:
@@
The approximation scheme of the potential will now be detailed. First, we write the dimension-full potential as $U(\phi_{2})=k\phi_{2}^{2}/2$, where $k$ serves as the spring constant of the dimension-full potential. The goal is to write the potential in the dimensionless form: $U^{\prime}(\varphi_{2})=k^{\prime}\varphi_{2}^{2}/2$, with $k^{\prime}$ being the spring constant of the dimensionless potential. This can be done by setting $k_{c}=U_{c}/2\phi_{c}^{2}$, in order to obtain $k^{\prime}=U^{\prime}/(\varphi_{2})^{2}$. Now, the particle’s motion is written as:
@@
which is analogous to the differential equation modelling a damped harmonic oscillator that oscillates at frequency $\omega^{2}=k^{\prime}\theta$ under damping coefficient $\lambda=\Lambda/2$. We deploy the ansatz: $\varphi_{2}(t^{\prime})=A\mathrm{e}^{\alpha t^{\prime}}$ in which $\alpha=-\lambda\pm\sqrt{\lambda^{2}-\omega^{2}}=-\lambda\pm\Omega$. We will focus on two possible situations which involve $\Omega^{2}>0$ or $\Omega^{2}<0$. Subsequently, these will be categorized as overdamped and underdamped, respectively.
@@ A.0.1 Overdamped
When $\Omega^{2}>0$, the solution to Eq. (7) is [41]:
@@
which describes an overdamped oscillator with initial conditions $A$ and $B$. To understand them, we first detail the consequences of the particle-like representation of the distribution of microstates. The average position of this distribution localizes to a position at any given time—represented by a particle in this case—while its average initial velocity will be zero. With this, the initial conditions are (i) $\varphi_{2}(t^{\prime}=0)=\varphi_{2}^{o}$ and (ii) $\mathrm{d}\varphi_{2}(t^{\prime}=0)/\mathrm{d}t^{\prime}=0$. Solving for these two initial conditions, and substituting the results back into Eq. (8), gives:
@@
Numerically solving Eq (9) for a final time $t^{\prime}=t_{f}^{\prime}$ and multiplying the result by $\sqrt{LC}$, yields $t_{\mathrm{CE}}=t_{f}^{\prime}\sqrt{LC}$.
@@ A.0.2 Underdamped
If $\Omega^{2}<0$, we must define $\widetilde{\omega}=\sqrt{\omega^{2}-\lambda^{2}}$ in order for the general solution to be written as [41]:
@@
Equation (10) describes an underdamped harmonic oscillator for which the coefficient $D$ and phase $\theta$ are solved using initial conditions. Subsequently, solving for them in the same manner as before and substituting the results into Eq. (10) produces:
@@ Appendix B Dwell Time
First, $\omega_{p,l}$ is the plasma frequency in the $l$th flux direction. It serves as the characteristic frequency of oscillations that are parallel to the escape direction; i.e., how frequently the particle approaches the barrier. To find it, we consider the dimension-full potential $U(\phi_{1},\phi_{2})$ and identify that we can perform a Taylor expansion in the $l$th coordinate around a minimum $(\phi_{1}=\phi_{1}^{*},\phi_{2}=\phi_{2}^{*})$. Of particular interest, and up to a constant, we look to the following second order terms of the potential:
$U(\phi_{1},\phi_{2})\approx\dfrac{1}{2}\dfrac{\partial^{2}U}{\partial\phi_{1}^{2}}\bigg{|}_{(\phi_{1}^{*},\phi_{2}^{*})}(\phi_{1}-\phi_{1}^{*})^{2}~{},\$ (12)
as well as:
$U(\phi_{1},\phi_{2})\approx\dfrac{1}{2}\dfrac{\partial^{2}U}{\partial\phi_{2}^{2}}\bigg{|}_{(\phi_{1}^{*},\phi_{2}^{*})}(\phi_{2}-\phi_{2}^{*})^{2}~{}.$ (13)
We identify the second partial derivative taken with respect to $\phi_{l}$ evaluated at the minimum to be the spring constant $k_{l}$. From here, we take the corresponding plasma frequency to be $\omega_{p,l}=\sqrt{k_{l}/m}$ for which the dimension-full mass $m=C$.
Furthermore, $\Delta U_{b}$ indicates the barrier height between a local minimum and a local maximum, $k_{B}$ is the Boltzmann constant, and $T$ is the circuit’s operating temperature. The escape rate is exponentially damped by the energy barrier because the thermal environment guarantees particles are exponentially unliklely to have energies that exceed the thermal energy scale. $\Gamma_{l}$ characterizes how many escape events are expected out of the minima per second. Finally, the dwell time is the reciprocal of Eq. (11): $t_{\mathrm{d},l}=1/\Gamma_{l}$. Explicitly, the plasma frequencies of interest for the dimensionless barriers $\Delta U_{01}^{\prime}$ ($\Delta U_{11}^{\prime}$ and $\Delta U_{10}^{\prime}$) correspond to the flux direction $l=1$ ($l=2$). We utilize the minimum dwell time to calculate the robustness $t_{\mathrm{d}}/t_{\mathrm{CE}}$.
@@ Appendix C Control Erasures of Interest
Figure 8 exhibits all control erasure protocols of interest. Appropriate successive executions of a subset of these protocols lead to performing a NAND gate. Note that this is not an exhaustive illustration of possible control erasure protocols, but only represents those of current interest. Table 2 shows the indices, signs and relationships of the relevant circuit parameters for a particular CE. The nonzero barrier-control parameter is determined by $i$. Meanwhile, the magnitudes of the tilt parameters that offset the bias and barrier changes to the potential are related by $|\varphi_{j\mathrm{x}}|>|\varphi_{i\mathrm{x}}|$, respectively. By using the effective magnitudes of the circuit parameters detailed in Sec. V.3 and successively executing effective CE protocols, a robust NAND gate can be performed.
Figure 8: All control erasure protocols of interest here, displayed with arrow-based notation first shown in Fig. 3.
@@
Table 2: Tabulated relationships between experimental circuit parameters—including their indices and signs—and the CE notation introduced in Sec. III. +()+\;(-)+ ( - ) corresponds to a positive (negative) sign of a particular circuit parameter.